OpenStudy (anonymous):

7. Show all work. Mary purchased a package of 13 different plants, but she only needed 9 plants for planting. In how many ways can she select the 9 plants from the package to be planted?

OpenStudy (anonymous):

your job is to compute "13 choose 9" written as \(\dbinom{13}{9}\) and computed via \[\dbinom{13}{9}=\frac{13\times 12\times 11\times 10}{4\times 3\times 2}\] cancel first, multiply last or use a calculator http://www.wolframalpha.com/input/?i=13+choose+9

OpenStudy (anonymous):

thanks i will go and visit that website